NSSYNC

version 1.1.91, 1 December 2014

Sergey Poznyakoff

Published by the Free Software Foundation, 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301 USA Copyright \begin{matrix}c \\ ◯ \end{matrix} 2012 Sergey Poznyakoff Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, an no specific Front- or Back-Cover texts.

Short Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Configuration File. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 4 Nssync Configuration . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5 Invocation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6 Exit Codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7 How to Report a Bug . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 A GNU Free Documentation License. . . . . . . . . . . . . . . . . . . . . . . 23 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Configuration File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3.1 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.2 Pragmatic Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 3.3 Statements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.4 Preprocessor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Nssync Configuration . . . . . . . . . . . . . . . . . . . . . . . . . 11

4.1 General Settings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 SQL Access . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 4.3 Synchronization Block. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

5 Invocation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6 Exit Codes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 7 How to Report a Bug . . . . . . . . . . . . . . . . . . . . . . . . . 21 Appendix A GNU Free Documentation License

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33

1 Introduction

BIND, the most frequently used DNS server, normally keeps its zone data in zone files. This approach becomes inconvenient when the number of zones grows beyond a certain limit. When this happens, the obvious solution is to move all data to a database and make named read it from there. Recent versions of BIND include dynamically loadable zones (DLZ) feature^{1}, which makes it possible to use such databases directly. However, DLZ has problems of its own, one of them being that it is unable to propagate glue records^{2}.

The nssync utility provides an alternative solution, which makes it possible to keep your zone data in an SQL^{3} database without using DLZ and with glue records working.

It does so by periodically polling the database to determine which data have changed recently and converting the database into BIND zone files.

1 See http://bind-dlz.sourceforge.net/.

2 See: http://permalink.gmane.org/gmane.network.dns.bind9.dlz/2078, http://

blog.gmane.org/gmane.network.dns.bind9.dlz/month=20110101.

3 As of version 1.1.91 only MySQL is supported.

2 Overview

The nssync utility is normally started periodically from crontab. Upon startup it reads its configuration file, which supplies the necessary program settings. Then, if the settings require so, it verifies that no other copy of the nssync is already running. Further on, it parses the named configuration file named.conf to determine several settings needed for its further operation, in particular, the value of the ‘directory’ statement in the ‘options’ block.

Once these preliminary operations are over, nssync starts its main task. Its configuration file defines, among other data, one or more synchronization blocks. Each such block defines SQL statements which return information about DNS zones as well as the location of named configuration file where the zone statements for these zones are to be stored (it is supposed that this file is included somewhere in the main named.conf file). For each synchronization block, the utility retrieves the zone data from the database and formats them into separate zone files. Each of these files is then compared to an already existing one (locations of the zone files are defined in the synchronization block they pertain to). If the files differ, new zone file replaces the old one and a flag is set indicating that the named daemon needs to be restarted in order to read new configuration.

When this stage is finished, nssync reloads the name server (if required) and exits.

Several command line options can be supplied in order to modify the program’s behavior. In particular, it is possible to check the configuration file syntax or even instruct the utility to do everything, except modifying the zone files (a so-called dry-run mode). This allows you to debug your configuration before actually starting using nssync.

3 Configuration File

Nssync reads its settings from a configuration file nssync.conf located normally in the system configuration directory (usually /etc or /usr/local/etc, depending on compile-time options).

This chapter describes the syntax of that file in general. The chapter that follows describes the nssync-specific settings in detail.

The configuration file consists of statements and comments.

There are three classes of lexical tokens: keywords, values, and separators. Blanks, tabs, newlines and comments, collectively called white space are ignored except as they serve to separate tokens. Some white space is required to separate otherwise adjacent keywords and values.

3.1 Comments

Comments may appear anywhere where white space may appear in the configuration file. There are two kinds of comments: single-line and multi-line comments. Single-line comments start with ‘#’ or ‘//’ and continue to the end of the line:

     # This is a comment
     // This too is a comment

Multi-line or C-style comments start with the two characters ‘/*’ (slash, star) and continue until the first occurrence of ‘*/’ (star, slash).

Multi-line comments cannot be nested. However, single-line comments may well appear within multi-line ones.

3.2 Pragmatic Comments

Pragmatic comments are similar to usual single-line comments, except that they cause some changes in the way the configuration is parsed. Pragmatic comments begin with a ‘#’ sign and end with the next physical newline character.

#include <file>
#include file

Include the contents of the file file. If file is an absolute file name, both forms are equivalent. Otherwise, the form with angle brackets searches for the file in the include search path, while the second one looks for it in the current working directory first, and, if not found there, in the include search path. The default include search path is:

           1. prefix/share/nssync/1.1.91/include
           2. prefix/share/nssync/include

where prefix is the installation prefix.

#include_once <file>
#include_once file

Same as #include, except that, if the file has already been included, it will not be included again.

#line num
#line num "file"

This line causes the parser to believe, for purposes of error diagnostics, that the line number of the next source line is given by num and the current input file is named by file. If the latter is absent, the remembered file name does not change.

# num "file"

This is a special form of #line statement, understood for compatibility with the c preprocessor.

In fact, these statements provide a rudimentary preprocessing features. For more sophisticated ways to modify configuration before parsing, see Section 3.4 [Preprocessor], page 8.

3.3 Statements

A simple statement consists of a keyword and value separated by any amount of whitespace. Simple statement is terminated with a semicolon (‘;’).

The following is a simple statement:

     standalone yes;
     pidfile /var/run/slb.pid;

A keyword begins with a letter and may contain letters, decimal digits, underscores (‘_’) and dashes (‘-’). Examples of keywords are: ‘expression’,

‘output-file’.

A value can be one of the following:

number A number is a sequence of decimal digits.

boolean A boolean value is one of the following: ‘yes’, ‘true’, ‘t’ or ‘1’, meaning true, and ‘no’, ‘false’, ‘nil’, ‘0’ meaning false.

unquoted string An unquoted string may contain letters, digits, and any of the following characters: ‘_’, ‘-’, ‘.’, ‘/’, ‘@’, ‘*’, ‘:’.

quoted string A quoted string is any sequence of characters enclosed in doublequotes (‘"’). A backslash appearing within a quoted string introduces an escape sequence, which is replaced with a single character according to the following rules:

Sequence  Replaced with
\a  Audible bell character (ASCII 7)
\b  Backspace character (ASCII 8)
\f  Form-feed character (ASCII 12)
\n  Newline character (ASCII 10)

\r

Carriage return character (ASCII 13)

\t

Horizontal tabulation character

(ASCII 9)

\v

Vertical tabulation character

(ASCII 11)

\\  A single backslash (‘\’)
\"  A double-quote.

Table 3.1: Backslash escapes In addition, the sequence ‘\newline’ is removed from the string. This allows to split long strings over several physical lines, e.g.:

     "a long string may be\
      split over several lines"

If the character following a backslash is not one of those specified above, the backslash is ignored and a warning is issued. Two or more adjacent quoted strings are concatenated, which gives another way to split long strings over several lines to improve readability. The following fragment produces the same result as the example above:

     "a long string may be"
     " split over several lines"

Here-document A here-document is a special construct that allows to introduce strings of text containing embedded newlines. The <<word construct instructs the parser to read all the following lines up to the line containing only word, with possible trailing blanks. Any lines thus read are concatenated together into a single string. For example:

     <<EOT A multiline string EOT

The body of a here-document is interpreted the same way as a double-quoted string, unless word is preceded by a backslash (e.g. ‘<<\EOT’) or enclosed in double-quotes, in which case the text is read as is, without interpretation of escape sequences. If word is prefixed with - (a dash), then all leading tab characters are stripped from input lines and the line containing word. Furthermore, if - is followed by a single space, all leading whitespace is stripped from them. This allows to indent here-documents in a natural fashion. For example:

     <<- TEXT
         The leading whitespace will be
         ignored when reading these lines.
     TEXT

It is important that the terminating delimiter be the only token on its line. The only exception to this rule is allowed if a heredocument appears as the last element of a statement. In this case a semicolon can be placed on the same line with its terminating delimiter, as in:

     help-text <<-EOT
             A sample help text.
     EOT;

list A list is a comma-separated list of values. Lists are enclosed in parentheses. The following example shows a statement whose value is a list of strings:

     alias (test,null);

In any case where a list is appropriate, a single value is allowed without being a member of a list: it is equivalent to a list with a single member. This means that, e.g.

     alias test;

is equivalent to

     alias (test);

A block statement introduces a logical group of statements. It consists of a keyword, followed by an optional value, and a sequence of statements enclosed in curly braces, as shown in the example below:

server srv1 {
host 10.0.0.1;
community "foo";
}

The closing curly brace may be followed by a semicolon, although this is not required.

3.4 Preprocessor

Before actual parsing, the configuration file is preprocessed. The built-in preprocessor handles only file inclusion and #line statements (see Section 3.2 [Pragmatic Comments], page 5), while the rest of traditional preprocessing facilities, such as macro expansion, is supported via m4, which serves as external preprocessor.

The detailed description of m4 facilities lies far beyond the scope of this document. You will find a complete user manual in Section “GNU M4” in GNU M4 macro processor. For the rest of this subsection we assume the reader is sufficiently acquainted with m4 macro processor.

The external preprocessor is invoked with -s flag, which instructs it to include line synchronization information in its output. This information is then used by the parser to display meaningful diagnostic.

An initial set of macro definitions is supplied by the pp-setup file, located

in prefix/share/nssync/1.1.91/include directory.

The default pp-setup file renames all m4 built-in macro names so they all start with the prefix ‘m4_’. This is similar to GNU m4 --prefix-builtin option, but has an advantage that it works with non-GNU m4 implementations as well.

4 Nssync Configuration

4.1 General Settings

These settings modify the behavior of nssync as a whole.

[Configuration] pidfile file

At startup, check if file already exists and is owned by an existing process. Exit if so. Use this statement to avoid accidentally running two copies of nssync simultaneously.

[Configuration] tempdir dir

Sets the name for the temporary directory. This is a directory where nssync creates temporary zone files. The argument must point to an existing directory.

[Configuration] check-ns bool

If set to true, nssync will check the list of NS servers prior to creating a zone file. The file will be created only if IPv4 address of one of the servers matches one of the IP addresses of the host on which nssync is run.

[Configuration] named-conf file

Defines the full pathname of the named configuration file. Default is

/etc/named.conf.
[Configuration] bind-include-path list

Sets include search path for include directives found in BIND configuration. The argument is either a single directory or a list of directories (see Section 3.3 [Statements], page 6).

[Configuration] zonefile-pattern pat

Defines the pattern for zone file names. The name of each zone file is created by expanding variable references in the pat argument. The following variable references are defined:

$zone ${zone} Name of the zone, without the trailing dot.

$synctag ${synctag}

Zone synchronization tag (see Section 4.3 [Synchronization Block], page 13).

Both notations (with and without braces) are equivalent. The notation with curly braces should be used if the reference is immediately followed by a letter. The default zone file pattern is ‘$zone.$synctag’.

[Configuration] zone-conf pat

Defines the pattern for zone configuration file, i.e. a file containing zone statements. The handling of pat is similar to that in zonefile-pattern, except that only the ‘$synctag’ reference is defined.

[Configuration] compare-command cmd

Defines a command to be used for comparing two zone files. The cmd must be a command taking two files as its arguments and returning 0 if they are the same or non-zero if they differ. Nssync uses this command to determine whether a particular zone has changed. The following variable references are expanded in cmd:

$oldfile ${oldfile} Old zone file.

$newfile

${newfile} New zone file.

The default compare-command value is:

    cmp $oldfile $newfile > /dev/null
[Configuration] reload-command cmd

Defines a command to reload the nameserver. The default is

‘/usr/sbin/rndc reload’.

4.2 SQL Access

The following statements define the database server and the database to use:

[Configuration] host hostname[:port-or-socket]

Defines the SQL server IP and port. The hostname can be either the server IP address or its hostname. The port-or-socket part, if supplied, can be either the number of TCP port to use instead of the default 3306 or the full pathname of the UNIX socket. In the latter case hostname is effectively ignored.

[Configuration] database name

Sets the database name.

[Configuration] ssl-ca file

Defines the name of the Certificate Authority (CA) file.

There are two ways to supply database access credentials. The simplest one is by using user and password statements:

[Configuration] user name

Sets SQL user name.

[Configuration] password arg

Sets SQL user password.

The drawback of this approach is that the password appears in plaintext, which means the permissions of the nssync.conf file must be tightened so as to avoid its compromise.

The following two statements provide an alternative, more safe and flexible way of setting access credentials:

[Configuration] sql-config-file file

Read MySQL configuration from the option file file. See Section “optionfiles” in MySQL Manual, for a description of MySQL option file format.

[Configuration] sql-config-group name

Read the named group from the SQL configuration file.

To illustrate their use, suppose your nssync.conf file contains the following:

    sql-config-file /etc/nssync.my;
    sql-config-group nssync;

The the /etc/nssync.my will contain the actual SQL access configuration, which can look as in the example below:

    [nssync]
    socket = /var/db/mysql.sock
    database = dns
    user = root
    pass = guessme
[Configuration] slave-status-file file

Use this statement if nssync reads data from a slave database. It allows you to avoid recreating zone files if the database information has not changed since the previous run. If this statement is present, nssync will save the state of the SQL slave in file. Upon startup, it will read these data and compare them with the current state. If they are the same, it will exit immediately.

4.3 Synchronization Block

A synchronization block defines a set of zones to be synchronized from the database and configures SQL statements which return the zone data. This set is identified by synchronization tag, supplied as the argument to the sync statement:

# Define a synchronization block.

    sync tag {

# zone configuration file

      zone-conf pat;

# pattern for new zone file names

   zonefile-pattern pat;

# add these statements to each generated zone file

   add-statements text;

# a query for retrieving SOA records

   soa-query string;

# a query for retrieving NS and similar records

   ns-query string;

# a query for retrieving the rest of RRs

   rr-query string;

# a query for retrieving RRs from reverse delegation zones

   rev-rr-query string;
 }

Statements within the sync block configure the zones:

[Configuration] zone-conf pat

Defines the pattern for the name of zone configuration file for zones in this synchronization block. If not supplied, the global zone-conf statement will be used instead (see [zone-conf], page 12).

[Configuration] zonefile-pattern pat

Defines the pattern for zone file names. If not supplied, the global zonefile-pattern statement will be used instead (see [zonefile-pattern], page 11).

[Configuration] add-statements text

Append text to each generated zone statement. For example, the following can be used to redefine forwarders and query ACLs for zones in this synchronization block:

    add-statements <<EOT
      forwarders { /* empty */ };
      allow-query { local-query-only; };
    EOT;

Notice the use of the here-document construct.

The following statements define which zones pertain to this particular synchronization block:

[Configuration] soa-query string

A query for retrieving SOA records.

[Configuration] ns-query string

A query for retrieving NS and similar records. Use the ‘$zone’ reference for the zone name.

[Configuration] rr-query string

A query for retrieving the rest of RRs. Use the ‘$zone’ reference for the zone name.

[Configuration] rev-rr-query string

A query for retrieving RRs from reverse delegation zones. Use the ‘$zone’ reference for the zone name.

Here is an example of a working sync directive:

sync external {
zone-conf "/var/namedb/nssync/zones.external";
zonefile-pattern "/var/namedb/external/db.${zone}";
soa-query    "select zone, ttl, type, data, resp_person, "
"serial, refresh, retry, expire, minimum "
"from dns_soa where type='SOA' "
"and view='external' order by zone";
ns-query     "select ttl, type, data "
"from dns_soa where zone='$zone' "
"and type<>'SOA' and view='external'";
rr-query     "select host, ttl, type, mx_priority, "
"case when type='TXT' then "
"concat('\"', data, '\"') "
"else data end "
"from dns_records "
"where zone='$zone' and view='external' "
"order by 1";
rev-rr-query "select host, ttl, type, mx_priority, "
"case when type='TXT' then "
"concat('\"', data, '\"') "
"else data end "
"from dns_records "
"where zone='$zone' and view='external' "
"order by cast(host as unsigned)";
}

5 Invocation

The nssync is normally invoked periodically from a crontab, e.g.:

    */5 * * * *  /usr/sbin/nssync | \
      /usr/bin/logger -t nssync -p local1.err

The following table summarizes available command line options:

-E Preprocess configuration file and exit.

-c file
--config-file=file

Use file instead of the default configuration file.

-f

--force Proceed even if slave status has not changed (see [slave-status-file], page 13).

-n
--dry-run

Do nothing, print almost everything; implies --debug --stderr. Use additional --debug options to get even more info.

-t

--lint Parse configuration file and exit. The return status is 0 if the syntax is OK, and 78 if errors were detected (see Chapter 6 [Exit Codes], page 19).

-D symbol=value
--define=symbol[=value]

Define a preprocessor symbol.

-I dir
--include-directory=dir

Add include directory.

--no-preprocessor

Disable preprocessing.

--preprocessor=command

Use command instead of the default preprocessor.

-d

--debug Increase debug level.

-X
--debug-lexer

Debug configuration file lexer.

-x
--debug-parser

Debug configuration file parser.

--config-help

Show configuration file summary

-V
--version

Print program version.

-h

--help Give this help list. --usage Give a short usage message.

6 Exit Codes

Apart from issuing a descriptive error message, nssync attempts to indicate the reason of its termination by its error code. As usual, a zero exit code indicates normal termination. The table below summarizes all possible error codes. For each error code, it indicates its decimal value and its symbolic name from include/sysexits.h (if available).

0 EX OK Program terminated correctly.

64 EX USAGE

The program was invoked incorrectly, e.g. an invalid option was given, or an erroneous argument was supplied to an option.

69 EX UNAVAILABLE

The program exited due to some error not otherwise described in this table.

70

EX SOFTWARE Some internal software error occurred.

78 EX CONFIG

An error in the configuration file was detected.

7 How to Report a Bug

Email bug reports to gray+nssync@gnu.org.ua. Please include a detailed description of the bug and information about the conditions under which it occurs, so we can reproduce it. To facilitate the task, the following list shows the basic set of information that is needed in order to find the bug:

• Package version you use.

• A detailed description of the bug.

• Conditions under which the bug appears.

• It is often helpful to send the contents of config.log file along with your bug report. This file is created after running ./configure in the nssync source root directory.

Appendix A GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000, 2001, 2002, 2007, 2008 Free Software Foundation, Inc.

http://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law. A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language. A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them. The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none. The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words. A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”. Examples of suitable formats for Transparent copies include plain ascii without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only. The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text.

The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages.

If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct

from that of the Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified

Version, as the publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adja-

cent to the other copyright notices. F. Include, immediately after the copyright notices, a license notice

giving the public permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections

and required Cover Texts given in the Document’s license notice. H. Include an unaltered copy of this License.

I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included in the Modified Version.

N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in title with any Invariant Section.

O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties— for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work. In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document.

If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License. However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns. See http://www.gnu.

org/copyleft/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008. The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

       Copyright (C)  year  your name.
       Permission is granted to copy, distribute and/or modify this document
       under the terms of the GNU Free Documentation License, Version 1.3
       or any later version published by the Free Software Foundation;
       with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
       Texts.  A copy of the license is included in the section entitled ``GNU
       Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

         with the Invariant Sections being list their titles, with
         the Front-Cover Texts being list, and with the Back-Cover Texts
         being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Concept Index 33 34 Nssync

Concept Index

This is a general index of all issues discussed in this manual.

#

#include. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 #include_once . . . . . . . . . . . . . . . . . . . . . . . . . 5 #line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

A

add-statements. . . . . . . . . . . . . . . . . . . . . . . 14

B

bind-include-path . . . . . . . . . . . . . . . . . . . 11 block statement . . . . . . . . . . . . . . . . . . . . . . . . 8 boolean value . . . . . . . . . . . . . . . . . . . . . . . . . . 6

C

check-ns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 Comments in a configuration file. . . . . . . . 5 comments, pragmatic. . . . . . . . . . . . . . . . . . . 5 compare-command . . . . . . . . . . . . . . . . . . . . . 12 configuration file statements . . . . . . . . . . . . 6

D

database . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

E

escape sequence . . . . . . . . . . . . . . . . . . . . . . . . 6

H

here-document . . . . . . . . . . . . . . . . . . . . . . . . . 7 host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

L

list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

M

m4 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 multi-line comments . . . . . . . . . . . . . . . . . . . . 5

N

named-conf . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 ns-query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 nssync.conf . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

P

password . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 pidfile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 pp-setup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 pragmatic comments . . . . . . . . . . . . . . . . . . . 5 preprocessor. . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

Q

quoted string. . . . . . . . . . . . . . . . . . . . . . . . . . . 6

R

reload-command. . . . . . . . . . . . . . . . . . . . . . . 12 rev-rr-query . . . . . . . . . . . . . . . . . . . . . . . . . 15 rr-query . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

S

simple statements . . . . . . . . . . . . . . . . . . . . . . 6 single-line comments . . . . . . . . . . . . . . . . . . . 5 slave-status-file . . . . . . . . . . . . . . . . . . . 13 soa-query . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 sql-config-file . . . . . . . . . . . . . . . . . . . . . 13 sql-config-group . . . . . . . . . . . . . . . . . . . . 13 ssl-ca . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 statement, block . . . . . . . . . . . . . . . . . . . . . . . 8 statement, simple . . . . . . . . . . . . . . . . . . . . . . 6 statements, configuration file . . . . . . . . . . . 6 string, quoted . . . . . . . . . . . . . . . . . . . . . . . . . . 6 string, unquoted. . . . . . . . . . . . . . . . . . . . . . . . 6 sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 synchronization block . . . . . . . . . . . . . . . 3, 13 synchronization tag . . . . . . . . . . . . . . . . . . . 13

T

tempdir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

U

user . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Z

zone-conf. . . . . . . . . . . . . . . . . . . . . . . . . 12, 14 zonefile-pattern. . . . . . . . . . . . . . . . . 11, 14